# Cosmic Collider Gravitational Waves sourced by Right-handed Neutrino production from Bubbles: Testing Seesaw, Leptogenesis and Dark Matter

### v1

@@ III.3 GW From Bubble Collision
FOPTs proceed via the nucleation of bubbles, their expansion, collision and thermalization into light particles, and GWs are produced during this process Caprini_2016. In the transition process, some of the released energy goes into heating up the plasma, while the rest is mostly carried by the scalar field configuration (bubble wall) and/or the bulk motion of the surrounding fluid. Gravitational wave production by such localized structure of energy around the walls are studied in detail over the past decade in literature and recent developments can be found in Refs. Jinno_2017; Jinno_2019; Cutting_2018. For runaway bubble configurations, GWs are primarily sourced by bubble wall collisions, i.e. the scalar field and produced particle densities. The GW spectrum obtained from the study in Caprini_2024; ghoshal2025complementaryprobeswarpedextra is given as
@@
Here $g_{*}$ is the total number of degrees of freedom present in the plasma. The SGWB spectrum then exhibits a broken-power-law frequency shape. In particular, at $f\ll f_{p}$ it behaves as $\Omega_{GW}h^{2}\propto f^{2.3}$ and at $f\gg f_{p}$, it behaves as $\Omega_{GW}h^{2}\propto f^{-2.4}$  [^3] This is much steeper than what Ref Caprini_2020 considered, i.e. $\Omega_{GW}h^{2}\propto f^{-1}$.. [^4] With an expanding FW background with more recent developments in lattice simulation, the scaling has been found to be slightly different in Ref. Lewicki:2025hxg.
@@ III.4 GW From Particle Production
In sec-II we saw that bubbles produced during a FOPT can create a particle density when they collide. If the energy momentum tensor of these produced particles then have a transverse traceless component, they can also source GWs. Indeed this contribution can be estimated via calculating the UETC in Eq. (28) for the produced particle distribution. This would be different from the UETC of the Bubble collisions. Then substituting the UETC in Eq. (30) and performing the integral one can get the GW spectrum coming from particles. For the particle production mechanism, this contribution is equal to Inomata:2024rkt
@@
Eq. (35) is robust for $k\lesssim\beta$ and runs into problematic behavior for $k\gtrsim\beta$. This originates from the short-distance structure of the two-point correlation function of the source in configuration space: as two bubbles cannot nucleate at the same spatial point, the correlation function must vanish in the limit of zero separation. This leads to a negative power spectrum in Fourier space for some frequencies in large momenta regime. We consequently discard this unphysical region and focus on the large scale physics for $k\lesssim\beta$. The expression Eq. (35) remains reliable up to $k/\beta\lesssim 0.75$ Inomata:2024rkt. In our analysis, we therefore present the GW spectrum only within this domain and truncate it beyond this value, indicating the cutoff by a vertical line in Fig. 2. The shape of the spectrum for $k\gtrsim\beta$ is a subject of ongoing research and is beyond the scope of this work.
In Fig. 2, we show the power-law–integrated (PLI) sensitivity curves for various upcoming gravitational-wave (GW) missions, following Ref. Thrane:2013oya. The shaded regions indicate the projected sensitivity reaches of the respective experiments. These curves assume that the GW spectrum from a first-order phase transition can be approximated by a power law, $\Omega_{\rm GW}\propto f^{b}$, where $b$ is the spectral index. A theoretical prediction lying within the shaded region corresponds to a signal detectable with high signal-to-noise ratio (SNR).While alternative approaches, such as peaked-integrated sensitivity curves (PISCs), may better characterize certain phase transition–induced GW spectra Schmitz:2020syl, we show the PLI curves for illustration and perform a dedicated SNR analysis next.
@@
In the same diagram we compare the GW signals as in Eq. (31,34) with respect to the PLI curves. It has been shown in the literature that for super-horizon modes the GW power spectrum sourced by a FOPT exhibits a universal scaling, $\Omega_{\rm GW}h^{2}\propto f^{3}$ (see Cai_2020). In the deep-IR regime, i.e. at frequencies corresponding to timescales longer than a Hubble time—the spectrum therefore follows the $f^{3}$ scaling due to causality PhysRevD.79.083519 rather than the $f^{1}$ behavior discussed for sub-horizon particle-production sources. To account for this, we impose an IR cutoff at the Hubble frequency, $f_{IR}=H(T_{*})$, and adopt the $f^{3}$ scaling for $f<f_{\rm IR}$. We stress that this cutoff is not a sharp, model-independent boundary: the precise transition point in the IR must be checked numerically and can vary with the details of the phase transition. [^5] See Eq. (33) of Ref. Inomata:2024rkt for analytical expression. Note we impose a similar IR cutoff for GWs coming from bubble collision as well.
@@ III.5 Signal-to-Noise-Ratio (SNR)
where we use $h=0.67$ and an observation time of $\tau=4$ years for all GW detectors except the PTA and SKA based GW detectors for which we take $\tau=20$ years. A detection threshold of SNR $\geq 10$ is applied. It is important to note that this formula for SNR calculation is got under the weak signal approximation, where the GW signal is much smaller than the instrumental noise Allen:1997ad. While it can at certain times overestimate the true SNR for strong signals, we adopt it here for both weak and strong GW signals for simplicity.
@@ III.5.1 Gravitational Wave Detectors
  • Ground-based interferometers: These detectors, such as LIGO/VIRGO/KAGRA (LVK)LIGOScientific:2016aoc; LIGOScientific:2016sjg; LIGOScientific:2017bnn; LIGOScientific:2017vox; LIGOScientific:2017ycc; LIGOScientific:2017vwq; KAGRA:2013rdx, aLIGO/aVIRGO LIGOScientific:2014pky; VIRGO:2014yos; LIGOScientific:2019lzm, AION Badurina:2021rgt; Graham:2016plp; Graham:2017pmn; Badurina:2019hst, Einstein Telescope (ET) Punturo:2010zz; Hild:2010id, and Cosmic Explorer (CE) LIGOScientific:2016wof; Reitze:2019iox, use interferometric techniques on the Earth’s surface to detect gravitational waves.
  • Space-based interferometers: Space-based detectors like LISA Baker:2019nia, BBO Crowder:2005nr; Corbin:2005ny; Cutler:2009qv, DECIGO, U-DECIGO Seto:2001qf; Yagi:2011wg, AEDGE AEDGE:2019nxb; Badurina:2021rgt, and $\mu$-ARES Sesana:2019vho are designed to detect gravitational waves from space, offering different advantages over ground-based counterparts.
  • Recasts of star surveys: Surveys such as GAIA/THEIA Garcia-Bellido:2021zgu utilize astrometric data from stars to indirectly infer the presence of gravitational waves.
  • Pulsar timing arrays (PTA): PTA experiments like SKA Carilli:2004nx; Janssen:2014dka; Weltman:2018zrl, EPTA EPTA:2015qep; EPTA:2015gke, and NANOGRAV NANOGRAV:2018hou; Aggarwal:2018mgp; NANOGrav:2020bcs use precise timing measurements of such pulsars to measure gravitational wave signatures.
@@ III.7 Bounds from BBN and CMB on dark radiation
The gravitational wave energy density has to obey certain observational bounds, for instance, should be smaller than the bound on dark radiation parameterized as the number of relativistic neutrino species $\Delta N_{\rm eff}$ Luo:2020fdt; Maggiore:1999vm,
@@
We typically do not consider the frequency dependence and neglect it, for this bound and set $\Omega_{\rm GW}\leq 5.6\times 10^{-6}\Delta N_{\rm eff}$ for the GW spectra that we calculate. BBN puts a bound on $\Delta N_{\rm eff}^{\rm BBN}\simeq 0.4$ Cyburt:2015mya. Planck plus BAO observations set the bound $\Delta N_{\rm eff}^{\rm Plack+BAO}\simeq 0.28$ Planck:2018vyg. Future projected bounds from future experiments are $\Delta N_{\rm eff}^{\rm Proj.}=0.014$ for CMB-HD CMB-HD:2022bsz, $\Delta N_{\rm eff}^{\rm Proj.}=0.05$ for CMB-Bharat CMB_Bharat, $\Delta N_{\rm eff}^{\rm Proj.}=0.06$ for CMB Stage IV Abazajian:2019oqj and NASA’s PICO mission Alvarez:2019rhd, $\Delta N_{\rm eff}^{\rm Proj.}\lesssim 0.12$ for CORE CORE:2017oje, the South Pole Telescope SPT-3G:2014dbx and Simons observatory detectors SimonsObservatory:2018koc.
@@ IV.1 Non-thermal Production of Right-handed Neutrino
where $y_{\phi}$ denotes the Yukawa coupling to $\phi$. If the bubbles of background field $\phi$ achieve runaway behavior, releasing vacuum energy $\Delta V=\lambda v_{\phi}^{4}$ as discussed earlier, then the produced RHN number density from bubble collision, given by Eq. (12) is (see cataldi2024leptogenesisbubblecollisions)
@@
Here the Planck mass $M_{pl}=1.22\times 10^{19}$ GeV and $\gamma_{max}$ is the maximum value of the Lorentz factor achieved by the bubble wall. Note due to the scalar self coupling, $\phi$ particles would also be produced from the phase transition. These can later decay into the $N$ states and change the yield significantly. The number density of $\phi$ particles produced from the process $\phi^{*}\rightarrow\phi\phi$ is cataldi2024leptogenesisbubblecollisions
@@ IV.2 DM formation and its GW probe
As noted in the introduction, observations indicate the presence of a non-luminous matter component comprising roughly $26\%$ of the Universe’s energy density — also known as dark matter (DM) Zwicky:1933gu. Since the Standard Model lacks a viable DM candidate, various beyond-the-Standard-Model (BSM) scenarios have been proposed Cirelli:2024ssz in order to explain the DM. In this section, we explore whether the RHNs introduced above—when rendered stable—can account for the observed DM relic abundance.
@@
With this in mind, we observe in Fig. 4 that in both panels the red curves intersect the $\Omega_{\rm DM}h^{2}=0.12$ line. This shows that for sufficiently large masses, $M_{N}\geq 10^{9}$ GeV, and couplings, $y_{\phi}\geq 0.01$, the RHN relic abundance matches the observed DM density. For small values of $M_{N}$, we have $M_{N}\lesssim y_{\phi}v_{\phi}$, and hence the mass essentially becomes constant ($\sim y_{\phi}v_{\phi}$) for $M_{N}\lesssim 1$ GeV. This leads to the nearly constant behavior of $\Omega_{\rm DM}h^{2}$ for $M_{N}\lesssim 1$ GeV, as seen in the left panel of Fig. 4. However, several assumptions entering Eq. (47) break down in the regime $M_{N}\lesssim v_{\phi}$, and results in this region should not be considered reliable [^10] If the particle gains mass mainly from the FOPT, i.e. $m\sim v_{\phi}$, its Compton wavelength $(\sim m^{-1})$ can exceed the thickness of the highly boosted bubble walls $(l_{w})$ at collisions. In this regime, the particle probes both vacua on either side of the bubble wall simultaneously, and the notion of a well-defined particle state during the collision breaks down, see Ref. Mansour:2023fwj for a detailed discussion. . Motivated by these considerations, we now perform a parameter scan in the $M_{N}-y_{\phi}$ plane.
@@
We see in Fig. 6, The GW signals coming from the FOPT event that produces correct relic abundance for the stable RHN dark matter can be successfully detected in multiple current and upcoming detectors. For mass as low as $M_{N}=10^{6}$ GeV, GW signal can be detected in the BBO detector for both particle production and bubble collision mechanisms whereas ET can detect only bubble collision GWs in this region. Part of the parameter space is already ruled out by LVK-$\mathcal{O}(3)$ data. On the other hand, for $M_{N}=10^{10}$ GeV, the corresponding GW signals coming from the bubble collisions can be detected in LISA and GW signals coming from particle production can be detected in SKA. Similarly, for very high mass $M_{N}=10^{14}$ GeV, the GW signals for both particle production and bubble collision can be detected in SKA. Note as was discussed in sec III and can be seen in Eq. (31&34), the corresponding GW signals coming from particle production and bubble collision has different scaling properties with frequency. This is particularly prominent in the low frequency regions of a given GW spectrum, and hence probing both contributions in separate detectors for the same PT parameters re-enforces the validity of the prediction.
@@ IV.3 Non-thermal Leptogenesis from Bubble collisions
As noted in the introduction, the observed baryon asymmetry and the origin of neutrino masses remain among the most pressing open questions in particle physics, and both can be addressed via leptogenesis. In this section, we explore the possibility of realizing leptogenesis through the production of non-thermal RHNs during a FOPT in this model  [^11] Leptogenesis in the context of supercooled phase transition has been studied in the literature but in a totally different Dasgupta:2022isg; Ghosh:2025non; Athron:2025pog.. We also investigate the corresponding gravitational-wave signals arising from both bubble-collision and particle-production mechanisms, analogous to the analysis performed for dark matter in the previous section.

### v2

@@ III.3 GW From Bubble Collision
FOPTs proceed via the nucleation of bubbles, their expansion, collision and thermalization into light particles, and GWs are produced during this process Caprini et al. (2016). In the transition process, some of the released energy goes into heating up the plasma, while the rest is mostly carried by the scalar field configuration (bubble wall) and/or the bulk motion of the surrounding fluid. Gravitational wave production by such localized structure of energy around the walls are studied in detail over the past decade in literature and recent developments can be found in Refs. Jinno and Takimoto (2017a, 2019a); Cutting et al. (2018a). For runaway bubble configurations, GWs are primarily sourced by bubble wall collisions, i.e. the scalar field and produced particle densities. The GW spectrum obtained from the study in Caprini et al. (2024); Ghoshal et al. (2025) is given as
@@
Here $g_{*}$ is the total number of degrees of freedom present in the plasma. The SGWB spectrum then exhibits a broken-power-law frequency shape. In particular [^3] This is much steeper than what Ref Caprini et al. (2020) considered, i.e. $\Omega_{GW}h^{2}\propto f^{-1}$. [^4] With an expanding FRW background with more recent developments in lattice simulation, the scaling has been found to be slightly different in Ref. Lewicki and Vaskonen (2025)., at $f\ll f_{p}$ it behaves as $\Omega_{GW}h^{2}\propto f^{2.3}$ and at $f\gg f_{p}$, it behaves as $\Omega_{GW}h^{2}\propto f^{-2.4}$.
@@ III.4 GW From Particle Production
In sec-II we saw that bubbles produced during a FOPT can create a particle density when they collide. If the energy momentum tensor of these produced particles then have a transverse traceless component, they can also source GWs. Indeed this contribution can be estimated via calculating the UETC in Eq. (28) for the produced particle distribution. This would be different from the UETC of the Bubble collisions. Then substituting the UETC in Eq. (30) and performing the integral one can get the GW spectrum coming from particles. For the particle production mechanism, this contribution is equal to Inomata et al. (2025)
@@
Eq. (35) is robust for $k\lesssim\beta$ and runs into problematic behavior for $k\gtrsim\beta$. This originates from the short-distance structure of the two-point correlation function of the source in configuration space: as two bubbles cannot nucleate at the same spatial point, the correlation function must vanish in the limit of zero separation. This leads to a negative power spectrum in Fourier space for some frequencies in large momenta regime. We consequently discard this unphysical region and focus on the large scale physics for $k\lesssim\beta$. The expression Eq. (35) remains reliable up to $k/\beta\lesssim 0.75$ Inomata et al. (2025). In our analysis, we therefore present the GW spectrum only within this domain and truncate it beyond this value, indicating the cutoff by a vertical line in Fig. 2. The shape of the spectrum for $k\gtrsim\beta$ is a subject of ongoing research and is beyond the scope of this work.
In Fig. 2, we show the power-law–integrated (PLI) sensitivity curves for various upcoming gravitational-wave (GW) missions, following Ref. Thrane and Romano (2013). The shaded regions indicate the projected sensitivity reaches of the respective experiments. These curves assume that the GW spectrum from a first-order phase transition can be approximated by a power law, $\Omega_{\rm GW}\propto f^{b}$, where $b$ is the spectral index. A theoretical prediction lying within the shaded region corresponds to a signal detectable with high signal-to-noise ratio (SNR).While alternative approaches, such as peaked-integrated sensitivity curves (PISCs), may better characterize certain phase transition–induced GW spectra Schmitz (2021), we show the PLI curves for illustration and perform a dedicated SNR analysis next.
@@
In the same diagram we compare the GW signals as in Eq. (31,34) with respect to the PLI curves. It has been shown in the literature that for super-horizon modes the GW power spectrum sourced by a FOPT exhibits a universal scaling, $\Omega_{\rm GW}h^{2}\propto f^{3}$ (see Cai et al. (2020)). In the deep-IR regime, i.e. at frequencies corresponding to timescales longer than a Hubble time—the spectrum therefore follows the $f^{3}$ scaling due to causality Caprini et al. (2009) rather than the $f^{1}$ behavior discussed for sub-horizon particle-production sources. To account for this, we impose an IR cutoff at the Hubble frequency, $f_{IR}=H(T_{*})$, and adopt the $f^{3}$ scaling for $f<f_{\rm IR}$. We stress that this cutoff is not a sharp, model-independent boundary: the precise transition point in the IR must be checked numerically and can vary with the details of the phase transition. [^5] See Eq. (33) of Ref. Inomata et al. (2025) for analytical expression. Note we impose a similar IR cutoff for GWs coming from bubble collision as well.
@@ III.5 Signal-to-Noise-Ratio (SNR)
where we use $h=0.67$ and an observation time of $\tau=4$ years for all GW detectors except the PTA and SKA based GW detectors for which we take $\tau=20$ years. A detection threshold of SNR $\geq 10$ is applied. It is important to note that this formula for SNR calculation is got under the weak signal approximation, where the GW signal is much smaller than the instrumental noise Allen and Romano (1999). While it can at certain times overestimate the true SNR for strong signals, we adopt it here for both weak and strong GW signals for simplicity.
@@ III.5.1 Gravitational Wave Detectors
  • Ground-based interferometers: These detectors, such as LIGO/VIRGO/KAGRA (LVK)Abbott and others (2016b, a, 2017e, 2017a, 2017b, 2017c, c), aLIGO/aVIRGO Aasi and others (2015); Acernese and others (2015); Abbott and others (2021), AION Badurina et al. (2021); Graham et al. (2016, 2017); Badurina and others (2020), Einstein Telescope (ET) Punturo and others (2010); Hild and others (2011), and Cosmic Explorer (CE) Abbott and others (2017d); Reitze and others (2019), use interferometric techniques on the Earth’s surface to detect gravitational waves.
  • Space-based interferometers: Space-based detectors like LISA Baker and others (2019), BBO Crowder and Cornish (2005); Corbin and Cornish (2006); Cutler and Holz (2009), DECIGO, U-DECIGO Seto et al. (2001); Yagi and Seto (2011), AEDGE El-Neaj and others (2020); Badurina et al. (2021), and $\mu$-ARES Sesana and others (2021) are designed to detect gravitational waves from space, offering different advantages over ground-based counterparts.
  • Recasts of star surveys: Surveys such as GAIA/THEIA Garcia-Bellido et al. (2021) utilize astrometric data from stars to indirectly infer the presence of gravitational waves.
  • Pulsar timing arrays (PTA): PTA experiments like SKA Carilli and Rawlings (2004); Janssen and others (2015); Weltman and others (2020), EPTA Lentati and others (2015); Babak and others (2016), and NANOGRAV Arzoumanian and others (2018); Aggarwal and others (2019); Arzoumanian and others (2020) use precise timing measurements of such pulsars to measure gravitational wave signatures.
@@ III.7 Bounds from BBN and CMB on dark radiation
The gravitational wave energy density has to obey certain observational bounds, for instance, should be smaller than the bound on dark radiation parameterized as the number of relativistic neutrino species $\Delta N_{\rm eff}$ Luo et al. (2021); Maggiore (2000),
@@
We typically do not consider the frequency dependence and neglect it, for this bound and set $\Omega_{\rm GW}\leq 5.6\times 10^{-6}\Delta N_{\rm eff}$ for the GW spectra that we calculate. BBN puts a bound on $\Delta N_{\rm eff}^{\rm BBN}\simeq 0.4$ Cyburt et al. (2016). Planck plus BAO observations set the bound $\Delta N_{\rm eff}^{\rm Plack+BAO}\simeq 0.28$ Aghanim and others (2020b). Future projected bounds from future experiments are $\Delta N_{\rm eff}^{\rm Proj.}=0.014$ for CMB-HD Aiola and others (2022), $\Delta N_{\rm eff}^{\rm Proj.}=0.05$ for CMB-Bharat 79, $\Delta N_{\rm eff}^{\rm Proj.}=0.06$ for CMB Stage IV Abazajian and Heeck (2019) and NASA’s PICO mission Alvarez and others (2019), $\Delta N_{\rm eff}^{\rm Proj.}\lesssim 0.12$ for CORE Delabrouille and others (2018), the South Pole Telescope Benson and others (2014) and Simons observatory detectors Ade and others (2019).
@@ IV.1 Non-thermal Production of Right-handed Neutrino
where $y_{\phi}$ denotes the Yukawa coupling to $\phi$. If the bubbles of background field $\phi$ achieve runaway behavior, releasing vacuum energy $\Delta V=\lambda v_{\phi}^{4}$ as discussed earlier, then the produced RHN number density from bubble collision, given by Eq. (12) is (see Cataldi and Shakya (2024))
@@
Here the Planck mass $M_{pl}=1.22\times 10^{19}$ GeV and $\gamma_{max}$ is the maximum value of the Lorentz factor achieved by the bubble wall. Note due to the scalar self coupling, $\phi$ particles would also be produced from the phase transition. These can later decay into the $N$ states and change the yield significantly. The number density of $\phi$ particles produced from the process $\phi^{*}\rightarrow\phi\phi$ is Cataldi and Shakya (2024)
@@ IV.2 DM formation and its GW probe
As noted in the introduction, observations indicate the presence of a non-luminous matter component comprising roughly $26\%$ of the Universe’s energy density — also known as dark matter (DM) Zwicky (1933). Since the Standard Model lacks a viable DM candidate, various beyond-the-Standard-Model (BSM) scenarios have been proposed Cirelli et al. (2024) in order to explain the DM. In this section, we explore whether the RHNs introduced above—when rendered stable—can account for the observed DM relic abundance.
@@
With this in mind, we observe in Fig. 4 that in both panels the red curves intersect the $\Omega_{\rm DM}h^{2}=0.12$ line. This shows that for sufficiently large masses, $M_{N}\geq 10^{9}$ GeV, and couplings, $y_{\phi}\geq 0.01$, the RHN relic abundance matches the observed DM density. For small values of $M_{N}$, we have $M_{N}\lesssim y_{\phi}v_{\phi}$, and hence the mass essentially becomes constant ($\sim y_{\phi}v_{\phi}$) for $M_{N}\lesssim 1$ GeV. This leads to the nearly constant behavior of $\Omega_{\rm DM}h^{2}$ for $M_{N}\lesssim 1$ GeV, as seen in the left panel of Fig. 4. However, several assumptions entering Eq. (47) break down in the regime $M_{N}\lesssim v_{\phi}$, and results in this region should not be considered reliable [^10] If the particle gains mass mainly from the FOPT, i.e. $m\sim v_{\phi}$, its Compton wavelength $(\sim m^{-1})$ can exceed the thickness of the highly boosted bubble walls $(l_{w})$ at collisions. In this regime, the particle probes both vacua on either side of the bubble wall simultaneously, and the notion of a well-defined particle state during the collision breaks down, see Ref. Mansour and Shakya (2025) for a detailed discussion. . Motivated by these considerations, we now perform a parameter scan in the $M_{N}-y_{\phi}$ plane.
@@
We see in Fig. 6, The GW signals coming from the FOPT event that produces correct relic abundance for the stable RHN dark matter can be successfully detected in multiple current and upcoming detectors. For mass as low as $M_{N}=10^{6}$ GeV, GW signal can be detected in the BBO and future LVK detectors for both particle production and bubble collision mechanisms whereas ET can detect only bubble collision GWs in this region. Part of the parameter space is already ruled out by LVK-$\mathcal{O}(3)$ data and further parameter space can be probed with future LVK observations. On the other hand, for $M_{N}=10^{10}$ GeV, the corresponding GW signals coming from the bubble collisions can be detected in LISA and GW signals coming from particle production can be detected in SKA. Similarly, for very high mass $M_{N}=10^{14}$ GeV, the GW signals for both particle production and bubble collision can be detected in SKA. Note as was discussed in sec III and can be seen in Eq. (31&34), the corresponding GW signals coming from particle production and bubble collision has different scaling properties with frequency. This is particularly prominent in the low frequency regions of a given GW spectrum, and hence probing both contributions in separate detectors for the same PT parameters re-enforces the validity of the prediction.
@@ IV.3 Non-thermal Leptogenesis from Bubble collisions
As noted in the introduction, the observed baryon asymmetry and the origin of neutrino masses remain among the most pressing open questions in particle physics, and both can be addressed via leptogenesis. In this section, we explore the possibility of realizing leptogenesis through the production of non-thermal RHNs during a FOPT in this model  [^11] Leptogenesis in the context of strong first order phase transition has been studied in the literature but in a different context Dasgupta et al. (2022); Ghosh et al. (2025); Athron et al. (2025); Ai et al. (2025); Cataldi et al. (2026).. We also investigate the corresponding gravitational-wave signals arising from both bubble-collision and particle-production mechanisms, analogous to the analysis performed for dark matter in the previous section.